tana-tanb=xandcotb-cota=y then prove that cot(a-b)=1/x+1/y

Grade:11

1 Answers

Prsahant Parashar
8 years ago
1/cot a – 1/cot b=x
cot b – cot a/cot a cot b=x
y/cot a cot b=x
y/x=cot a cot b
 
   cot(a-b)=cota cotb+1/cotb-cota
             =y/x+1/y
             =(y+x)/x/y
            =1/x+1/y